Use the definition of limit to prove that both of the sequences {1/ √ n } and { (−1)^n/ √ n } converges to 0.


1
Expert's answer
2021-01-26T11:55:37-0500

Solution 1:

We know that "\{a_n\}" converges to L, or, "\lim_{n \rightarrow \infty}=L" if "\forall" "\epsilon" ">0" "\exist" N"\in" "Z^+" such that "\forall" "n>N",

"|a_n-L|<\epsilon" .

Consider "|\dfrac{1}{\sqrt n}-0|=|\dfrac{1}{\sqrt n}|=\dfrac{1}{\sqrt n}<\epsilon"

"\Rightarrow 1<\epsilon\cdot\sqrt n \Rightarrow \dfrac{1}{\epsilon}<\sqrt n \Rightarrow n>\dfrac{1}{\epsilon^2}"

Proof: Let "\epsilon" ">0" . Choose N">\dfrac{1}{\epsilon^2}" . Then, "\forall" "n>N", "|\dfrac{1}{\sqrt n}-0|=|\dfrac{1}{\sqrt n}|"

Also, "n>N>\dfrac{1}{\epsilon^2} \Rightarrow \epsilon^2>\dfrac1 n \Rightarrow \sqrt{\dfrac1n}<\epsilon \Rightarrow {\dfrac1{\sqrt n}}<\epsilon"

Thus, "|\dfrac{1}{\sqrt n}-0|=\dfrac{1}{\sqrt n}<\epsilon"

Hence, by above statement, "\{\dfrac{1}{\sqrt n}\}" converges to 0.

Solution 2:

Consider "|\dfrac{(-1)^n}{\sqrt n}-0|=|\dfrac{(-1)^n}{\sqrt n}|=\dfrac{1}{\sqrt n}<\epsilon"

"\Rightarrow 1<\epsilon\cdot\sqrt n \Rightarrow \dfrac{1}{\epsilon}<\sqrt n \Rightarrow n>\dfrac{1}{\epsilon^2}"

Proof: Let "\epsilon" ">0" . Choose N">\dfrac{1}{\epsilon^2}" . Then, "\forall" "n>N", "|\dfrac{(-1)^n}{\sqrt n}-0|=|\dfrac{(-1)^n}{\sqrt n}|=\dfrac{1}{\sqrt n}"

Also, "n>N>\dfrac{1}{\epsilon^2} \Rightarrow \epsilon^2>\dfrac1 n \Rightarrow \sqrt{\dfrac1n}<\epsilon \Rightarrow {\dfrac1{\sqrt n}}<\epsilon"

Thus, "|\dfrac{(-1)^n}{\sqrt n}-0|=|\dfrac{(-1)^n}{\sqrt n}|=\dfrac{1}{\sqrt n}<\epsilon"

Hence, by above statement, "\{\dfrac{(-1)^n}{\sqrt n}\}" converges to 0.